OpenStudy (anonymous):

In triangle ABC, angle C is a right triangle, what is the measure of x? A. 42.1 B. 64.8 C. 23.1 D. 25.2

OpenStudy (anonymous):



OpenStudy (mathstudent55):

I'm sure you mean angle C is a right _angle_

OpenStudy (mathstudent55):

You have the lengths of al 3 sides of the triangle, so you can choose which trig function to use to get the angle measure.

OpenStudy (mathstudent55):

Keep in mind that for angle B, 28.3 is the adjacent leg 60.1 is the opposite leg 66.4 is the hypotenuse

OpenStudy (anonymous):

ya that is what i meant sorry

OpenStudy (mathstudent55):

Do you know the trig functions, sine, cosine and tangent?

OpenStudy (anonymous):

no im stuck with what do to do now, lol

OpenStudy (mathstudent55):

Have you ever learned SOH CAH TOA?

OpenStudy (mathstudent55):

It's a way to memorize the trig functions in terms of the ratios they are equal to.

OpenStudy (mathstudent55):

SOH means Sin = Opp/Hyp CAH means Cos = Adj/Hyp TOA means Tan = Opp/Adj

OpenStudy (anonymous):

ok, thanks i completely forgot how to use sohcahtoa

OpenStudy (mathstudent55):

Since you have all the lengths, you know what adj, opp and hyp are equal to. You can use any of the three trig functions to get the angle. sin x = 60.1/66.4 cos x = 28.3/66.4 tan x = 60.1/28.3

OpenStudy (anonymous):

alright thanks!

OpenStudy (mathstudent55):

sin x = opp/hyp sin x = 60.1/66.4, etc.

OpenStudy (mathstudent55):

wlcm